OpenStudy (anonymous):

which polynomial identity will prove that 37=64-27? difference squares difference of cubes sum of cubes square of a binomial

OpenStudy (shrutipande9):

if u r confused...d best way is to apply each option....64 is a square of what? or cube of what?

OpenStudy (solomonzelman):

\[37=4^3-3^3\]

OpenStudy (solomonzelman):

I would therefore think B.

OpenStudy (solomonzelman):

Guys makes sense?

OpenStudy (anonymous):

yes i think it does

OpenStudy (shrutipande9):

are u getting it @helpineedit

OpenStudy (solomonzelman):

OK, other questions then?

OpenStudy (anonymous):

what about for (3+6)^2=81 what would that be?

OpenStudy (solomonzelman):

\[ (3+6)^2=81\]\[ (3+6)^2=9^2\]you guesses?

OpenStudy (anonymous):

sum of cubes?

OpenStudy (solomonzelman):

cubes, where do you see cubes?

OpenStudy (solomonzelman):

I see squares, no?

OpenStudy (anonymous):

difference in squares then sorry

OpenStudy (solomonzelman):

Difference is subtracting, and you are not subtracting.

OpenStudy (shrutipande9):

see u r talking about polynomial identities...tell me what all identities u know?

OpenStudy (anonymous):

square of a binomial

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

and ik difference squares difference of cubes sum of cubes square of a binomial @shrutipande9

OpenStudy (solomonzelman):

The square of a binomial is always a trinomial.

OpenStudy (anonymous):

it has to be the sum of cubes then

OpenStudy (solomonzelman):

Binomial is something like 4x+16 it involves a variable and your question does not. Difference of squares is just (a+b)(a-b)=a^2-b^2 Sum of cubes. a^3 + b^3 = (a + b)(a^2 – ab + b^2) Difference of cubes. a^3 – b^3 = (a – b)(a^2 + ab + b^2)

OpenStudy (solomonzelman):

\[(3+6)^2=81\] doesn't really look like any of those.

OpenStudy (anonymous):

i know but it has to be i just dont know which one i though sum of cubes but i guess i was worng

hartnn (hartnn):

square of a binomial, 3^2 + 2*3*6 +6*2 = 81 the best among choices

hartnn (hartnn):

6^2 ***

OpenStudy (solomonzelman):

Yeah, it is far far from the truth though. It is invalid.

OpenStudy (solomonzelman):

(3+6) is not a binomial, and will never be it.

hartnn (hartnn):

the question just asks for ploynomial identity which will prove that.

OpenStudy (solomonzelman):

Well, it's not called a square of a binomial, is it?

OpenStudy (solomonzelman):

I agree it's the best choice, but the choices are all very bad.

hartnn (hartnn):

(3+6)^2 using the polynomial identity of square of binomila = 3^2 + 2*3*6 +6^2 = 81 hence proved.

OpenStudy (solomonzelman):

I can see this somewhat.

hartnn (hartnn):

:)

OpenStudy (solomonzelman):

I would not do that type of problem out of principle, if I god forbid had it.